OpenStudy (anonymous):

Show that if n is a positive integer then, [{1,1}{0,1}] = [{1,n}{0,1}] (these are supposed to be 2x2 matrices)

OpenStudy (anonymous):

oh, that was supposed to be [{1,1}{0,1}]^n = [{1,n}{0,1}]

OpenStudy (amistre64):

does the equal sign refer to row equivalence?

OpenStudy (amistre64):

oh, exponent stuff. if show means prove, then you might have to use induction

OpenStudy (amistre64):

establish a basis: 1 1 ^1 = 1 1 0 1 0 1 is true let n=k to define it in a more generic manner 1 1 ^k = 1 k 0 1 0 1 is assume to be true multiply both sides by 1 1 0 1 to get to the k+1 exponent 1 1 ^(k+1) = 1 k * 1 1 0 1 0 1 0 1 1 1 ^(k+1) = 1 k+1 0 1 0 0+1

OpenStudy (amistre64):

so if it is true for some integer k, it is true for some integer k+1 we know it is true for k=n=1

OpenStudy (anonymous):

thank you so much! I was thinking that I would have to use induction, but was not quite sure how to do it using matrices.

OpenStudy (amistre64):

i just went with a gut feeling :)